# Towards a Theory of Pragmatic Information

### v5

@@ 3.3 Pragmatic Independence of Messages
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})=\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime})$ (4)
@@ 4 Pragmatic Information as “Free Information,” By Way of Mutual Information
Pragmatic information can be interpreted as a kind of information analog of the Hemholtz free energy of statistical mechanics in the following sense: the Hemholtz free energy, $F$, is given by the formula
$F=U-TS.$
It is is a measure of the amount of energy in a system available to do useful work at a constant temperature, $T$, as opposed to the total energy, $U$, of the system ($S$ is the system’s entropy). Similarly, as we now show, pragmatic information is what remains when information irrelevant to making the decision is removed.
@@
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
@@
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $=\sum_{i,m}\varphi_{m}p_{i|m}\log\left[\frac{p_{i|m}}{q_{i}}\right]$
@@
$=\mathcal{I}(\mathcal{M};\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q}),$
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
@@
When $\mathbf{p}=\mathbf{q}$, the information needed to specify outcome $\omega_{i}\in\Omega$, i.e. the amount of information used in making a presumably useful decision, is the total amount of information transmitted by receipt of $m\in\mathcal{M}$, reduced by $H(\mathcal{M}|\Omega)$, the amount of information that is unneeded to specify the given outcome, and thus useless for making the decision. This difference is similar in spirit to free energy, the difference between the total energy and the internal energy, i.e. the information that is not available to do useful work. We can therefore characterize pragmatic information as “free information”; in fact, it is a generalization of the free information defined in [1].
@@
5 Pragmatic Utility and the Efficient Market Hypothesis
As indicated in the Introduction, a given message/message ensemble can, for a given decision maker, $\Delta$, have positive pragmatic information, yet misinform $\Delta$. This follows from our definition of pragmatic information as information that changes $\Delta$’s views of the state of the world, regardless of whether the new view is any more accurate, better, etc. than some base state. Nor is $\Delta$ guaranteed to have enough context to assign a value to that decision and thus to the messages leading to that decision. For example, what should be the value of a message that can prompt a single, terrible, but low probability outcome, as opposed to a message that leads to a set of more likely, but merely bad outcomes? Or can we be sure that we can assign values $V(m_{1}),V(m_{2}),V(m_{3})$ to messages $m_{1},m_{2},m_{3}\in\mathcal{M}$, such that $V(m_{1})<V(m_{2})$ and $V(m_{2})<V(m_{3})$ implies $V(m_{1})<V(m_{3})$?
Instead, we make the weaker assumption that we know enough about the implications of messages to classify them as being either
@@ 5 Pragmatic Utility and the Efficient Market Hypothesis
  • pragmatic disinformation (or simply “disinformation”) with respect to $\Delta$,
  • pragmatically useful with respect to $\Delta$.
We might, for example, make the above classification based on whether a given message is less useful, the same as, or more useful to $\Delta$ than not receiving the message at all. In any case, (2) is the same for pragmatic disinformation and pragmatically useful information, because that formula captures only the degree to which the probabilties of $\Delta$’s decisions, and not the value of these decisions to $\Delta$, have changed.
@@
  • $\mathcal{I}_{\Delta}$ is the set of pragmatically irrelevant messages with respect to $\Delta$.
  • $\mathcal{D}_{\Delta}$ is the set of pagmatically disinformative messages with respect to $\Delta$.
@@
The corresponding amounts of pragmatically irrelevant, disinformative, and useful messages can then be defined, as $\Phi_{\Delta}(\mathcal{I}_{\Delta};\Omega),\Phi_{\Delta}(\mathcal{D}_{\Delta};\Omega)$, and $\Phi_{\Delta}(\mathcal{U}_{\Delta};\Omega)$, respectively, with
@@
We now apply the above definitions to make a useful reformulation of the efficient market hypothesis, a cornerstone of financial economics (See, for example, [18].). According to this hypothesis, current asset prices, such as stock and bond prices, reflect all available information, so there is nothing to be gained by poring over charts of price histories or over annual reports.
@@
Clearly, the above characterization of market efficiency depends on the participant, as well as the message ensemble being considered. The market can therefore be efficient to some participants, but not others because information may be pragmatically useful to one market participant but not others [^7] This is certainly the case, for example, regarding knowledge of short term mispricings that can be exploited by automated trading systems, but not human traders.. That the market can be efficient with respect to certain message ensembles, but not others is already a staple of the relevant literature. [18], for example, distinguishes between the “weak”, “semi-strong”, and “strong” forms of the hypothesis that, respectively, price histories, all available public information, and all available information, including insider information, are irrelevant in predicting future asset price fluctuations. Nevertheless, our conceptual framework seems to provide one immediate insight: namely that a market is necessarily efficient with respect to a given participant and a given message ensemble if that participant is unable to extract any pragmatically useful information from that ensemble. One such situation is when the amount of incoming market information overwhelms the computational capacities of the participant, as the present author has found from personal experience!
@@ 6 Directions for Future Research: $\Delta$’s Computational Capacity and Beyond
Furthermore, the details of the computational apparatus that $\Delta$ is assumed to possess are relevant, not just for our efficient markets application, but also for biological evolution. Recall, per [3] and references therein, that the quasi-species is a simplfied model of abstract “replicators”, subject to Darwinian “survival of the fittest” in the presence of mutation. It is the interaction of the quasi-species with its environment that “decides” which replicator is, in fact, the fittest. Therefore, this interaction is the $\Delta$ of the present theory, because it is this interaction that “decides” the unknown identity of the fastest growing replicator. The identity of this replicator is the random variable, $\omega$, and the environmental details that determine growth and mutation rates are the “messages” that the quasi-species environment system processes. The whole point of the formulation of the quasi-species theory [19] was to demonstrate the existence of an “error catastrophe”, which is a limit to the information that the quasi-species “computation” could reliably accrue in any given situation.

### v6

@@ 3.3 Pragmatic Independence of Messages
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})=\Phi_{\Delta}({\mathcal{M}};\Omega)+\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime}).$ (4)
@@ 4 Pragmatic Information as “Free Information,” By Way of Mutual Information
Pragmatic information can be interpreted as a kind of information analog of the Hemholtz free energy of statistical mechanics [^5] See textbooks such as [20] for an introduction to this quantity., a measure of the amount of energy in a system available to do useful work. The Hemholtz free energy, $F$, is defined by the formula
$F=U-TS,$
where $U$ is the total energy of the system, $T$ is the temperature, and $S$ is the system’s entropy. Similarly, as we now show, pragmatic information is a measure of the information that is useful in making a decision.
@@
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q}).$
@@
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $=\sum_{i,m}p_{i,m}\log\left[\frac{p_{i|m}}{q_{i}}\right]$
@@
$=\mathcal{I}(\mathcal{M};\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q}),$
@@
When $\mathbf{p}=\mathbf{q}$, the information needed to specify outcome $\omega_{i}\in\Omega$, i.e. the amount of information used in making the decision, is the total amount of information transmitted by receipt of $m\in\mathcal{M}$, reduced by $H(\mathcal{M}|\Omega)$, the amount of information that is unneeded to specify the given outcome, and thus useless for making the decision. It is in this sense that pragmatic information is similar in spirit to free energy, the difference between the total energy and the internal energy, i.e. the information that is not available to do useful work. We can therefore characterize pragmatic information as “free information”; in fact, it is a generalization of the free information defined in [1]. If, in addition, there is one-to-one relationship between messages and outcomes, $H(\mathcal{M}|\Omega)=0$, and $\Phi(\mathcal{M};\Omega)=\mathcal{H}$ because all of the information transmitted is used in making the decision.
@@
5 On Disinformation and the Vexed Question of the Value of Information
The careful reader will note that we have made no attempt thus far to assign a value to the information in our incoming messages, however valuable they may be to the receiver. By avoiding any such mention of the value of information, we implicitly make the point that the value of a message is distinct from the amount of pragmatic information it contains (This point is obvious if we remind ourselves that pragmatic information is measured in bits, whereas value would be measured in dollars or other quantity that would prove useful to a decision maker.).
While the most obvious reason for our reluctance to assign such values is that the value of information is highly context dependent, there are several other difficulties with such an assignment. First, as indicated above, a given message/message ensemble can, for a given decision maker, $\Delta$, have positive pragmatic information, yet misinform $\Delta$. This follows from our definition of pragmatic information as information that changes $\Delta$’s views of the state of the world, regardless of whether the new view is any more accurate, better, etc. than some base state. Nor are we guaranteed to have enough context to assign a value to a decision and thus to the messages leading to that decision. For example, what should be the value of a message that might prompt a terrible outcome with low probability, but is more likely to prompt more beneficial outcomes? And how should that value compare to the value of another message that leads uniformly to outcomes that are bad, but not that bad? Even more problematic is our inability to guarantee an assignment of values $V(m_{1}),V(m_{2}),V(m_{3})$ to messages $m_{1},m_{2},m_{3}\in\mathcal{M}$, such that $V(m_{1})<V(m_{2})$ and $V(m_{2})<V(m_{3})$ implies $V(m_{1})<V(m_{3})$.
Nevertheless, the weaker assumption that messages are either
@@ 5 On Disinformation and the Vexed Question of the Value of Information
  • pragmatic disinformation (or simply “disinformation”) with respect to $\Delta$, or
  • pragmatically useful with respect to $\Delta$
is justified by [19]. These authors envision a situation in which information about the receiver’s environment may or may not promote the receiver’s viability. While the discussion in [19] uses biological examples, their framework is equally relevant for the receipt of environmental information that promotes/degrades a receiver’s economic viability. More broadly, we might make the above classification based on whether a given message is less useful, the same as, or more useful to $\Delta$ than not receiving the message at all. In any case, (2) is the same for pragmatic disinformation and pragmatically useful information, because that formula captures only the degree to which the probabilties of $\Delta$’s decisions, and not the value of these decisions to $\Delta$, have changed.
@@
  • $\mathcal{I}_{\Delta}$ is the set of pragmatically irrelevant messages with respect to $\Delta$,
  • $\mathcal{D}_{\Delta}$ is the set of pagmatically disinformative messages with respect to $\Delta$,
@@
The corresponding amounts of pragmatically irrelevant, disinformative, and useful messages can then be defined as $\Phi_{\Delta}(\mathcal{I}_{\Delta};\Omega),\Phi_{\Delta}(\mathcal{D}_{\Delta};\Omega)$, and $\Phi_{\Delta}(\mathcal{U}_{\Delta};\Omega)$, respectively, with
@@
6 A Reformulation of the Efficient Market Hypothesis
We now apply the above definitions to make a useful reformulation of the efficient market hypothesis, a cornerstone of financial economics (see, for example, [16]). According to this hypothesis, current asset prices, such as stock and bond prices, reflect all available information, so there is nothing to be gained by poring over charts of price histories or over annual reports.
@@ 6 A Reformulation of the Efficient Market Hypothesis
Thus, under this formulation, a market can be efficient to some participants, but not others because information may be pragmatically useful to one market participant but not others [^8] This is certainly the case, for example, regarding knowledge of short term mispricings that can be exploited by automated trading systems, but not human traders.. That the market can be efficient with respect to certain message ensembles but not others is already a staple of the relevant literature. [16], for example, distinguishes between the “weak”, “semi-strong”, and “strong” forms of the hypothesis that, respectively, price histories, all available public information, and all available information, including insider information, are irrelevant in predicting future asset price fluctuations. Nevertheless, our conceptual framework seems to provide one immediate insight: namely, that a market is necessarily efficient with respect to a given participant and a given message ensemble if that participant is unable to extract any pragmatically useful information from that ensemble. One such situation is when the amount of incoming market information overwhelms the computational capacities of the participant, as the present author has found from personal experience!
@@ 7 Directions for Future Research: $\Delta$’s Computational Capacity and Beyond
Furthermore, the details of the computational apparatus that $\Delta$ is assumed to possess are relevant, not just for our efficient markets application, but also for biological evolution. Recall, per [3] and references therein, that the quasi-species is a simplfied model of abstract “replicators”, subject to Darwinian “survival of the fittest” in the presence of mutation. It is the interaction of the quasi-species with its environment that “decides” which replicator is, in fact, the fittest. Therefore, this interaction is the $\Delta$ of the present theory, because it is this interaction that “decides” the unknown identity of the fastest growing replicator. The identity of this replicator is the random variable, $\omega$, and the environmental details that determine growth and mutation rates are the “messages” that the quasi-species environment system processes. The whole point of the original formulation of the quasi-species theory [17] was to demonstrate the existence of an “error catastrophe”, which is a limit to the information that the quasi-species “computation” could reliably accrue in any given situation.